OpenStudy (anonymous):

Log derivatives: y = e^3ln(x+1)

OpenStudy (anonymous):

d/dx ln(u) = u' / u

OpenStudy (anonymous):

So d/dx ln(x + 1) = ?

OpenStudy (anonymous):

is there a log rule?

OpenStudy (anonymous):

Also, is ln(x+1) a part of exponent or just 3?

OpenStudy (anonymous):

ln(x+1) is a part of the exponent and 3 is the coefficient

OpenStudy (anonymous):

Oh I see. Remember that 3ln(x+ 1) = ln(x + 1)³ and e^lnx = x So what does e^(3ln(x+1)) equal to?

OpenStudy (anonymous):

so what happens to the e?

OpenStudy (anonymous):

because i did product rule for the 3ln(x+1)

OpenStudy (anonymous):

It was removed, since e^ln(x) = x

OpenStudy (anonymous):

Instead, use the properties of logarithm. We know that n·lnx = lnxⁿ and e^lnf(x) = f(x) So e^(n·lnf(x)) = xⁿ So what does e^(3ln(x+1)) equal to?

OpenStudy (anonymous):

3(X+1)^2

OpenStudy (anonymous):

We know that e^(3ln(x+1)) = (x+1)³ Applying chain rule: f(g(x)) = f'(g(x)) · g'(x) Let f(x) be x³ and g(x) be x + 1. So (x+1)³ = 3(x + 1)² · 1 = 3(x+1)² So, you are correct, sir. Is this clear?

OpenStudy (anonymous):

lol yes thank you!

OpenStudy (anonymous):

Glad I helped lol.